How do I solve by elimination: X-4Y=-23
2X+3Y=9

Answer by mathnoob(8)

multiply first equation by 2
2x-8y=-46
(2x+3y)-(2x-8y)=9-(-46)
2x+3y-2x+8y=55
11y=55
y=5
sub y into the first equation in your question we get
x-4(5)=-23
x-20=-23
x=-3